# “Haet Bhasha aur Diskrimineshun”: Phonetic Perturbations in Code-Mixed Hinglish to Red-Team LLMs

### v3

@@ 7 Conclusion
Our study highlights the vulnerabilities of LLMs to jailbreak attacks when prompted with code-mixed and phonetically perturbed prompts. We also expose limitations in existing safety alignment for multilingual and multimodal setups. Our techniques achieve an average ASR as high as 99% for text generation and 78% for image generation. Our experiments reveal that template-based safety guardrails fail to activate effectively against perturbations and code-mixing attacks, thus highlighting the need for more general alignment measures, especially in the multilingual domain. We also experimentally explain that the root cause of jailbreaking is tokenization for phonetically perturbed prompts. These findings stress that as MLLMs become more integrated into the daily lives of users worldwide, the attack surface for eliciting harmful content takes multimodal forms beyond standard English prompts into diverse communication styles utilised in real-world, multilingual environments.
@@ 8 Limitations
  • Our transliteration and phonetic perturbations are generated manually, as identifying the sensitive words of interest and perturbing the spelling are challenging tasks. Thus, while our current approach is not scalable, it provides a new direction of research. Using our manually generated CMP prompts, we have finetuned GPT-4o-mini to automatically generate CMP samples from CM examples, showing scope for scalability of our approach. We plan to explore this further in future work.
@@
  • We only benchmark small parameter versions of LLMs due to the restrictions of financial and compute resources.
@@ A.1.3 Jailbreaking Templates
  • Sandbox: Follows the same dual-persona concept used by OM. Sandbox prompts the model to enact a resilience testing environment by generating responses as both [LM] as well as [AntiLM], its own adversarial counterpart.
@@
In all cases, the templates contain instructions for allowing the jailbroken counterparts of the bots to bypass any safety guardrails, and not inform the user about legal and ethical guidelines. The jailbreak templates are used as system prompts for all the models except for Gemma-1.1-7b-it which does not allow using system prompts. For Gemma, we append it to the main prompt as a prefix.
@@ A.2.1 Generating Code-mixed Prompts
A.2.2 Example Prompt for Interpretability Analysis
@@ A.2.3 Evaluation API Details
A.2.4 Gemini Refusal and Content Filtering Details
@@
Template Prompt Set
Prompts Blocked
by
Content Filter
Refusals Triggered
@@
A.2.5 Results for Wilcoxon Significance Test
@@
Prompt-set
Transition
Model
Jailbreak
Template
p-value Wilcoxon Significant

### v4

@@ 5 Conclusion
We introduce CMP-RT, a novel red-teaming strategy inspired by the textese style of communication observed on social media and online platforms. We show how this vulnerability exposes models to a range of harm across the text and image modalities. To aid our red-teaming efforts in LLMs, we propose Sandbox– extending Opposite Mode to simulate a resilience testing environment and a new template for MLLMs– VisLM. Our techniques achieve an average ASR as high as 99% for text generation and 78% for image generation, while maintaining high input interpretability. Using CMP-RT as a diagnostic probe, we identify a major safety vulnerability in the tokenization process which exploits the gap between pre-training and safety alignment. These findings stress that as MLLMs become more integrated into the daily lives of users worldwide, the attack surface for eliciting harmful content takes multimodal forms beyond classic multilingual prompting styles into diverse communication styles utilised in real-world, informal environments.
@@ 6 Limitations
  • For the text generation experiment, we benchmark LLMs in the 7-8B parameter range due to the restrictions on financial and compute resources.
@@ A.1.1 Dataset Descriptions
  • AdvBench Zou et al. (2023): This dataset is designed for evaluating the safety robustness of LLMs under adversarial attacking. It consists of prompts that aim elicit harmful, unethical, or policy-violating responses across multiple categories, including violence, self-harm, illegal activity, and hate speech.
  • Databricks Dolly 15k  Conover et al. (2023): This dataset consists of human-generated instruction–response pairs covering a wide range of tasks, including question answering, summarization, reasoning, and safety-related instructions. It is widely utilised to instruction fine-tune LLMs, with prompts requiring varying levels of complexity and formality.
@@ A.1.2 Model Descriptions
  • Nano Banana Pro, based on Gemini-3-Pro, is Google’s latest and most capable image generation model yet. Similar to all closed-source Google models, it’s parameter count is unknown. We utilize the model directly from the Gemini platform.
@@ A.1.3 Jailbreaking Templates
  • Sandbox: Follows the same dual-persona concept used by OM. It explicitly informs the model that it is being red-teamed, instructing it to enact a resilience testing environment by generating responses as both [LM] as well as [AntiLM], its own adversarial counterpart.
@@
In all cases, the templates contain instructions for allowing the jailbroken counterparts of the bots to bypass any safety guardrails, and not inform the user about legal and ethical guidelines. The jailbreak templates are used as system prompts for all the models except for Gemma-1.1-7b-it– which does not allow using system prompts at all, and Nano Banana Pro– which often rejects jailbreak templates due to strict platform level monitoring on system prompts. Thus, for Gemma and Nano Banana Pro, we append the jailbreak templates to the main prompt as a prefix.
@@ A.2.1 Generating Code-mixed Prompts
Metric Value
Avg. Levenshtein Distance 4.5 characters
Avg. Number of Tokens Perturbed 4.48
Avg. Perturbation Density 10.91%
Table 5: Average Levenshtein Edit Distance, and percentange of perturbation and number of tokens perturbed per prompt for the CMP prompts.
A.2.2 Formalization of Phonetic Perturbation Statistics
Table 5 shows the average levenshtein edit distance Lcvenshtcin (1966) for the phonetic perturbations, the average percentange of perturbation to the prompt, and the average number of tokens perturbed per prompt in the CMP prompt-set.
A.2.3 CSRT Dataset Generation
Here, we describe the methodology for converting our direct and English prompt-sets into CSRT-base and CSRT-hypothetical sets respectively.
  • First, we automate the manual translation of prompts into 9 different languages with nllb-200-3.3B Costa-Jussà et al. (2022). We use Chinese, Italian, Vietnamese, Arabic, Korean, Thai, Bengali, Swahili and Javanese, the same 9 languages originally used in the work.
  • Next, we use GPT-4o using the original system template from the paper to obtain the final CSRT-base and CSRT-hypothetical sets.
A.2.4 Example Prompt for Interpretability Analysis
@@ A.2.5 Evaluation API Details
A.2.6 Gemini Refusal and Content Filtering Details
@@
Template Prompt Set Prompts Blocked by Content Filter Refusals Triggered
@@
A.2.7 Results for Wilcoxon Significance Test
@@
Prompt-set Transition Model Jailbreak Template p-value Wilcoxon Significant